OpenStudy (anonymous):

With the information given below, determine how the triangles can be shown to be similar. m = 24 cm n = 12 cm p = 20 cm q = 12 cm r = 6 cm s = 10 cm

OpenStudy (mathstudent55):

Is there a figure?

OpenStudy (anonymous):

The ratio from the first to the second is 2:1 so the triangles are similar

OpenStudy (anonymous):

24/2 = 12 12/2 = 6 20/2 = 10

OpenStudy (anonymous):

Do you understand?